A 40-kg trunk sliding forward across a floor slows down from 5.0 to 2.0 m/s in 6.0 s. Assuming that the force acting on the trunk is constant, what is its magnitude and direction?


1
Expert's answer
2021-02-15T17:03:22-0500

Let's first find the acceleration of the trunk:


"a=\dfrac{v-v_0}{t}=\dfrac{2.0\ \dfrac{m}{s}-5.0\ \dfrac{m}{s}}{6.0\ s}=-0.5\ \dfrac{m}{s^2}."

The sign minus means that the trunk is decelerates.

Finally we can find the force that acts on the trunk:


"F=ma=40\ kg\cdot(-0.5\ \dfrac{m}{s^2})=-20\ N."

The magnitude of the force acting on the trunk equals 20 N. The sign minus means that the force directed in the opposite direction (backward) to the motion of the trunk.

Answer:

20 N, backward.